# Guided Bayesian Optimization: Data-Efficient Controller Tuning with Digital Twin

### v1

@@ Abstract
This article presents the guided Bayesian optimization algorithm as an efficient data-driven method for iteratively tuning closed-loop controller parameters using an event-triggered digital twin of the system based on available closed-loop data. We define a controller tuning framework independent of the controller or the plant structure. Our proposed methodology is model-free, making it suitable for nonlinear and unmodelled plants with measurement noise. The objective function consists of performance metrics modeled by Gaussian processes. We utilize the available information in the closed-loop system to identify and progressively maintain a digital twin that guides the optimizer, improving the data efficiency of our method. Switching the digital twin on and off is triggered by data-driven criteria related to the digital twin’s uncertainty estimations in the BO tuning framework. Effectively, it replaces much of the exploration of the real system with exploration performed on the digital twin. We analyze the properties of our method in simulation and demonstrate its performance on two real closed-loop systems with different plant and controller structures. The experimental results show that our method requires fewer experiments on the physical plant than Bayesian optimization to find the optimal controller parameters.
@@ I Introduction
One can directly incorporate available plant data in the control parameter tuning and the control algorithm. Direct data-driven controller design techniques can relieve the modeling task of complicated systems. Data-driven inversion-based control [9] and reachable zone-based pure data-driven controller design using only noisy input-output data of a trajectory of the system [10] are all independent from the system’s model for controller tuning. Nevertheless, such direct methods are usually less data-efficient than their model-based counterparts, especially when the controller needs to learn about complex real systems with limited data [11].
Integrating the measurements with simulation behaviors allows the data-driven controller to adapt effectively to the real world [12]. For example, the iterative learning controller combines the process measurements with the optimization framework to provide robustness in the presence of repetitive disturbances and plant-model mismatch [13, 14, 15]. Recently, researchers have designed reinforcement learning-based controllers that over-explore a broad family of simulated models of a complicated system [16]. According to [17], over-exploring the simulations improves the gradient-based position control policy to train with fewer experiments on a real robot. Furthermore, one can also use the model-based policy iteration to ensure regional optimality by locally approximating the system’s behavior [18]. [19] explicitly estimates the dynamics model uncertainty to design a robust model-based reinforcement learning method that reduces the sample complexity of controlling a robot arm.
@@
Bayesian optimization (BO) is a data-driven technique that iteratively tunes the controller parameters in a limited number of experiments using a surrogate model of the cost function [23]. To increase the BO controller tuning efficiency, one can integrate all available prior information about the system into the optimizer [24]. For instance, [25] uses coarse physical knowledge about the Plasma Spray Process to encapsulate a linear dependence between the process inputs in the prior mean distribution. Furthermore, [26] uses BO to tune the parameters of a goal-oriented model predictive controller that sets the reference input to another cascaded control system to reduce the energy consumption of a thermal management system. This way, the predictive controller’s state-space model serves as prior knowledge for BO, ruling out the search space that contradicts the plant’s physics. In both examples, BO is guided by prior information about the process.
This work proposes Guided Bayesian Optimization (Guided BO) method that incorporates available closed-loop system information into the optimization. Our data-efficient method reduces the required real system operation to tune the industrial feedback controller parameters iteratively. In our approach, automated controller tuning seeks a global optimum of a cost function that estimates the desired closed-loop system behavior. We use the available data without additional system operation to create an approximate digital twin (DT) of the system. We only take measurements on the real system when the BO surrogate model has sufficiently low uncertainty. Otherwise, we use our digital twin to estimate the system performance and guide the optimizer. This way, we tune the controller parameters with fewer experiments on the real system, reducing the operating cost. We demonstrate the enhanced performance of the guided BO algorithm in simulation and two real industrial feedback control systems.
@@ II Data-Driven Controller Tuning Problem
We consider a feedback control system with possibly a nonlinear plant. We have no assumption on the parametric controller structure. Let $\theta\in\Theta$ be the controller parameter vector inside a feasible set $\Theta\subseteq\mathbb{R}^{n_{\theta}}$. We assume the feasible set $\Theta$ is known (see Sections V and VI).
Given the controller parameter vector $\theta$, we denote the cost function $\hat{J}(\theta)$, where $\hat{J}:\Theta\to\mathbb{R}$ is a scalar metric reflecting the closed-loop performance. We want to simultaneously optimize multiple performance metrics denoted by $f_{i}:\Theta\rightarrow\mathbb{R}$ for $i=1,...,n_{\mathrm{f}}$ and $n_{\mathrm{f}}>1$. These metrics are measured based on the tracking error signal $e$, the difference between the reference and output signal. While one approach would be to optimize all the metrics separately determining the Pareto frontier[27], we adopt the classical weighted aggregation method to convert our multi-objective problem to a single-objective problem [28]. So the overall performance cost function $\hat{J}(\theta)$ is defined as
@@ III-A Bayesian Optimization-based Controller Tuning
where $\text{K}_{n}$ is defined as $[k(\theta_{\xi_{1}},\theta_{\xi_{2}})]\bigr{\rvert}_{\xi_{1},\xi_{2}=1}^{n}\in\mathbb{R}^{n\times n}$, $\sigma_{\epsilon}$ is the standard deviation of the additive noise to the predictive performance uncertainty, $\mathbb{I}$ is the identity matrix, vector $\mu(\mathrm{X}_{n})=[\mu(\theta_{1}),...,\mu(\theta_{n})]^{\mathsf{T}}\in\mathbb{R}^{n}$, and the vector $k_{n}(\theta)\in\mathbb{R}^{n}$ is equal to $[k(\theta,\theta_{1}),...,k(\theta,\theta_{n})]^{\mathsf{T}}$.
@@
We exhaustively grid search for the next parameter $\theta_{n+1}$ within the feasible set $\Theta$ to solve the following optimization problem:
@@ III-B Digital Twin
In the BO-based controller tuning method, we already acquire informative data in each experiment so one could build a plant model. Starting from an initial training data set with size $N_{0}$, we build the digital-twin data set $\tilde{\mathcal{D}}$ at experiment $n$ as
@@
We estimate an approximate model of the real plant called DT plant denoted by $\tilde{\mathrm{G}}$ based on $\tilde{\mathcal{D}}$. Assume $\tilde{\mathrm{G}}$ is a linear time-invariant transfer function estimated by the refined instrumental variable method [39]. This DT plant model is computationally cheap and fast to evaluate. So our DT is the same closed-loop system but with a DT plant instead of the real plant.
@@ III-C Guided Bayesian Optimization
Algorithm 1 summarizes our guided BO method. At each BO iteration using the physical plant $\mathrm{G}$, we collect the control signal $u$ and the plant output $y$ to build DT data set $\tilde{\mathcal{D}}$ and identify a DT plant model $\tilde{\mathrm{G}}$. The DT plant is updated each time the performance of the closed-loop control system with the physical plant is measured. Based on the GP model uncertainty, the guided BO algorithm gains more information from the updated DT. More specifically, we activate the DT when two conditions
@@
are satisfied where $\eta_{1}\geq 0$ is the DT activation threshold, $v_{n}$ is the variance of the BO’s posterior GP model calculated at the next query parameter vector $\theta^{+}$, $e_{\text{RMSE}}$ is RMSE defined in (11), and $\eta_{2}$ is the RMSE threshold related to the real plant output signal to noise ratio (cf. Section V). When the DT is activated, we use the DT instead of the actual system to estimate the cost value denoted by $\tilde{\hat{J}}$ at $\tilde{\theta}:=\theta^{+}$. We add a Gaussian noise with zero mean and $\sigma_{\epsilon}^{2}$ variance to the estimated performance to obtain $\tilde{J}:=\tilde{\hat{J}}+\epsilon$ where $\epsilon\sim\ \mathcal{N}(0,\sigma_{\epsilon}^{2})$. This estimated performance $\tilde{J}$ is then added to the BO training data set in Line 11 of Algorithm 1. In Section V-C, we numerically study the DT activation threshold $\eta_{1}$.
We continuously monitor the DT quality to maintain optimum controller parameters when the real plant changes. Each time we measure the performance on the real system $J$, we compare it with its estimated value by DT denoted by $\tilde{J}$. Suppose that the system behavior has considerably changed. In that case, we remove all previous data from the DT dataset. More specifically, if the normalized difference between the measured and estimated costs is considerable, i.e., $\frac{|J-\tilde{J}|}{J}>\tilde{\delta}$ where $\tilde{\delta}$ is DT re-initialization threshold, then we reinitialize DT dataset with the recent measurements as in Line 24 of Algorithm 24. If the system significantly changes, a new safe boundary for the optimization parameters must be determined, as demonstrated in [40, 41], for example.
@@
1:Input: Set feasible set $\Theta$, initial control parameters $\Theta_{\text{init}}$, weight vector $\mathrm{w}$, DT sampling length $\tilde{n}$, DT activation parameters $\eta_{1}$ and $\eta_{2}$, DT stopping parameter $\eta_{3}$, DT re-initialization parameter $\tilde{\delta}$, and $\tilde{n}_{\text{EI}}$, and maximum experiments on $\mathrm{G}$
@@
3:while maximum experiments on $\mathrm{G}$ is not reached do
@@
7:     if DT activation criteria in (III-C) is met then
@@
11:              Update BO training data: $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{(\tilde{\theta},\tilde{J})\right\}$
@@
16:         Remove all DT data $(\tilde{\theta},\tilde{J})$ from $\mathcal{D}$ in 3
17:         Set $\theta^{+}=\theta_{m}\in\mathcal{D}\text{ s.t. }\forall\theta\in\mathcal{D}:J(\theta_{m})<J(\theta)$
@@
23:     if DT re-initialization criterion is met then
@@
An appropriate stopping criterion for the optimization process is necessary for the practical implementation of the algorithm. We terminate the BO after a fixed number of iterations on the real system. When using DT, we need a stopping condition that ensures prompt termination of the optimization. We use the stopping criterion that depends on the maximal expected improvement over previous iterations and the current expected improvement proposed in [29]. Once we have activated the DT plant to predict the closed-loop response, we stop the BO iterations on DT depending on the ratio between the last expected improvement and the maximal expected improvement realized so far. This way, we avoid unnecessary iterations using the DT plant where we anticipate no further improvement. Notably, we stop BO on DT at iteration $n$ when, for $\tilde{n}_{{\text{EI}}}$ consecutive iterations, we meet
$a_{\text{EI},\text{n}}(\theta)\leq\eta_{3}\>\underset{n-\tilde{n}_{EI}\leq j\leq n-1}{\mathrm{max}}\>a_{\text{{\text{EI}}},\text{j}}(\theta),$ (13)
where $0<\eta_{3}$ is a predefined threshold to stop BO iteration on DT as in [42]. We stop BO iteration on DT when the expected improvement does not significantly improve compared with the improvements in the last $\tilde{n}_{EI}$ iterations.
As soon as we stop using DT, we remove all the DT data $(\tilde{\theta},\tilde{J})$ from $\mathcal{D}$. Then, we update the query parameter vector $\theta^{+}$ with the optimum observed parameter vector in the BO training data set $\mathcal{D}$ as defined in line 17 of Algorithm 1. We eventually proceed with the actual system to measure the overall performance at $\theta^{+}$.
@@ IV-A System Structure and Identification
[…]
$\mathrm{G}(s):=\frac{L}{s^{2}+\tau_{1}s+\tau_{2}}\,\mathrm{e}^{-sl},$ (14)
where Table I summarizes its identified parameters.

### v2

@@ Abstract
This article presents the guided Bayesian optimization (BO) algorithm as an efficient data-driven method for iteratively tuning closed-loop controller parameters using a digital twin of the system. The digital twin is built using closed-loop data acquired during standard BO iterations, and activated when the uncertainty in the Gaussian Process model of the optimization objective on the real system is high. We define a controller tuning framework independent of the controller or the plant structure. Our proposed methodology is model-free, making it suitable for nonlinear and unmodelled plants with measurement noise. The objective function consists of performance metrics modeled by Gaussian processes. We utilize the available information in the closed-loop system to progressively maintain a digital twin that guides the optimizer, improving the data efficiency of our method. Switching the digital twin on and off is triggered by our data-driven criteria related to the digital twin’s uncertainty estimations in the BO tuning framework. Effectively, it replaces much of the exploration of the real system with exploration performed on the digital twin. We analyze the properties of our method in simulation and demonstrate its performance on two real closed-loop systems with different plant and controller structures. The experimental results show that our method requires fewer experiments on the physical plant than Bayesian optimization to find the optimal controller parameters.
@@ I Introduction
One can directly incorporate available plant data in the control parameter tuning and the control algorithm which can relieve the modeling task of complicated systems. Convolution-based data-driven method [9] simulates the feedback system without building any model of it that accounts for the controller nonlinearities. This method can control rotary actuators by accurately estimating the control input and system response[10], but it holds a strong assumption that the plant is a linear time-invariant system. Reachable zone-based pure data-driven controller design uses only noisy input-output data of a trajectory of the system [11] independent from the system’s model for controller tuning. Nevertheless, such direct methods are usually less data-efficient than their model-based counterparts, especially when the controller needs to learn about complex real systems with limited data [12].
Integrating the measurements with simulation behaviors allows the data-driven controller to adapt effectively to the real world [13]. Learning from measurement can improve the tracking accuracy of an inverse compensator for a fiber-based piezoelectric actuator [14]. Furthermore, the iterative learning controller combines the process measurements with the optimization framework to provide robustness in the presence of repetitive disturbances and plant-model mismatch [15, 16, 17]. A digital twin (DT) is a virtual replica of a real-world system that continuously updates with measurements from its physical counterpart. DT enables online monitoring and optimization of complex manufacturing processes. For example, [18] explains how a liquid-level control system can be monitored by a DT that encompasses virtual twins of all subsystems. These components are then synchronized with the physical counterpart through a proper bridge between the subsystems such that one can extend the application to control the real process variables [19, 20]. However, there is a gap between the ”twinning” property of DTs in industrial systems which limits the DT accuracy. To mitigate the system-twin gap in the controller tuning, [21] incorporates a generic learning algorithm into the DT that enables active braking control of a race car. [22] challenges the sample efficiency of the learning-based algorithms in comparison with the local gradient-based black-box optimization methods. [23] uses the model-based policy iteration to ensure regional optimality by locally approximating the system’s behavior.
@@
Bayesian optimization (BO) is a data-driven technique that iteratively tunes the controller parameters in a limited number of experiments using a surrogate model of the cost function [27]. In complex industrial applications, conducting each of the iterative experiments can involve time-consuming measurements [28]. To increase the BO controller tuning efficiency, one can integrate all available process information into the optimizer [29]. For instance, [30] uses coarse physical knowledge about the Plasma Spray Process to encapsulate a linear dependence between the process inputs in the prior mean distribution. In [31] the process information is incorporated in BO to save experimental time in a tedious additive manufacturing application. Furthermore, [32] uses BO to tune the parameters of a goal-oriented model predictive controller that sets the reference input to another cascaded control system to reduce the energy consumption of a thermal management system. This way, the predictive controller’s state-space model serves as prior knowledge for BO, ruling out the search space that contradicts the plant’s physics. In previous examples, BO is guided by prior information about the process.
This work proposes Guided Bayesian Optimization (Guided BO) method that incorporates available closed-loop system information into the optimization. Our data-efficient method reduces the required real system operation to tune the industrial feedback controller parameters iteratively. In our approach, automated controller tuning seeks a global optimum of a cost function that estimates the desired closed-loop system behavior. We use the available data without additional system operation to create a DT. We only take measurements on the real system when the BO surrogate model has sufficiently low uncertainty. Otherwise, we use our DT to estimate the system performance and guide the optimizer. This way, we tune the controller parameters with fewer experiments on the real system, reducing the operating cost. We demonstrate the enhanced performance of the guided BO algorithm in simulation and two real industrial feedback control systems.
@@ II Data-Driven Controller Tuning Problem
We consider a feedback control system with possibly a nonlinear plant. Fig. 1 visualizes the components of our controller tuning problem. We have no assumption on the parametric controller structure. Let $\theta\in\Theta$ be the controller parameter vector inside a feasible set $\Theta\subseteq\mathbb{R}^{n_{\theta}}$. We assume the feasible set $\Theta$ is known (see IV-B and VI-B). For the convergence properties of the closed-loop system, we choose the set $\Theta$ such that the step response of the system is bounded in a fixed period.
Given the controller parameter vector $\theta$, we denote the cost function $\hat{J}(\theta)$, where $\hat{J}:\Theta\to\mathbb{R}$ is a scalar metric reflecting the closed-loop performance. There is no assumption on the convexity of the cost function. So $\hat{J}(\theta)$ may have multiple local minima. We want to simultaneously optimize multiple performance metrics denoted by $f_{i}:\Theta\rightarrow\mathbb{R}$ for $i=1,...,n_{\mathrm{f}}$ and $n_{\mathrm{f}}>1$. These metrics are measured based on the tracking error signal $e$, the difference between the reference and output signal. While one approach would be to optimize all the metrics separately, determining the Pareto frontier[33], we adopt the classical weighted aggregation method to convert our multi-objective problem to a single-objective problem [34]. So the overall performance cost function $\hat{J}(\theta)$ is defined as
@@ III-A Bayesian Optimization-based Controller Tuning
where $\text{K}_{n}$ is defined as $k(\theta_{\xi_{1}},\theta_{\xi_{2}})]\bigr\rvert_{\xi_{1},\xi_{2}=1}^{n}\in\mathbb{R}^{n\times n}$, $\sigma_{\epsilon}$ is the standard deviation of the additive noise to the predictive performance uncertainty, $\mathbb{I}$ is the identity matrix, vector $\mu(\mathrm{X}_{n})=[\mu(\theta_{1}),...,\mu(\theta_{n})]^{\mathsf{T}}\in\mathbb{R}^{n}$, and the vector $k_{n}(\theta)\in\mathbb{R}^{n}$ is equal to $k(\theta,\theta_{1}),...,k(\theta,\theta_{n})]^{\mathsf{T}}$.
@@
To optimize the acquisition function for the next parameter $\theta_{n+1}$, we exhaustively grid search with a fixed number of data within the feasible set $\Theta$:
@@
The computed sequence of $\theta_{1},\theta_{2},\theta_{3},$... by this iterative optimization converges to the proper solution of (1) [45, 46].
The computational complexity of BO controller tuning at iteration $m$ is heavily influenced by inversion of the covariance matrix $\mathrm{K}_{n}$ in (6) and is approximated by $O(m.n^{3})$ where $n$ is the size of training data set $\mathcal{D}$.
@@ III-B Digital Twin
In the BO-based controller tuning method, we already acquire informative data in each experiment so one could constantly learn a digital counterpart of the closed loop system. Starting from an initial training data set with size $N_{0}$, we build the digital-twin data set $\tilde{\mathcal{D}}$ at experiment $n$ as
@@
We estimate an approximate model of the real plant called DT plant denoted by $\tilde{\mathrm{G}}$ based on $\tilde{\mathcal{D}}$. Assume $\tilde{\mathrm{G}}$ is a linear time-invariant transfer function estimated by the refined instrumental variable method [47]. This DT plant model is computationally cheap and does not require fast processing. So our DT is the same closed-loop system but with a DT plant instead of the real plant.
@@ III-C Guided Bayesian Optimization
Algorithm 1 summarizes our guided BO method. At each BO iteration using the physical plant $\mathrm{G}$ and given the query control parameters $\theta^{+}$, we collect the control signal $u$ and the plant output $y$ to build DT data set $\tilde{\mathcal{D}}$ and identify a DT plant model $\tilde{\mathrm{G}}$. The DT plant is updated each time the performance of the closed-loop control system with the physical plant is measured. Based on the GP model uncertainty, the guided BO algorithm gains more information from the updated DT. More specifically, we activate the DT when two conditions
@@
are satisfied where $\eta_{1}\geq 0$ is the DT activation threshold, $v_{n}$ is the variance of the BO’s posterior GP model calculated at the next query parameter vector $\theta^{+}$, $e_{\text{RMSE}}$ is RMSE defined in (11), and $\eta_{2}$ is the RMSE threshold and is chosen based on the noise standard deviation as $\eta_{2}:=3\sigma_{\epsilon}$. We estimate $\sigma_{\epsilon}^{2}$ by measuring the performance metrics in five experiments with identical controller gains in the feasible set. When the DT is activated, we use the DT instead of the actual system to estimate the cost value denoted by $\tilde{\hat{J}}$ at $\tilde{\theta}:=\theta^{+}$. We add a Gaussian noise with zero mean and $\sigma_{\epsilon}^{2}$ variance to the estimated performance to obtain $\tilde{J}:=\tilde{\hat{J}}+\epsilon$ where $\epsilon\sim\ \mathcal{N}(0,\sigma_{\epsilon}^{2})$. This estimated performance $\tilde{J}$ is then added to the BO training data set in Line 11 of Algorithm 1. In Section V-C, we numerically study the DT activation threshold $\eta_{1}$.
We continuously monitor the DT quality to maintain optimum controller parameters when the real plant changes. Each time we measure the performance on the real system $J$, we compare it with its estimated value by DT denoted by $\tilde{J}$. Suppose that the system behavior has considerably changed. In that case, we discard all previous data from the DT dataset. More specifically, if the normalized difference between the measured and estimated costs is considerable, i.e.,
$\frac{|J-\tilde{J}|}{J}>\tilde{\delta}$, (13)
where $\tilde{\delta}$ is DT re-initialization threshold, then we reinitialize DT dataset with the recent measurements as in Line 24 of Algorithm 24. If the system significantly changes, a new safe boundary for the optimization parameters must be determined, as demonstrated in [48, 49], for example.
@@
1:Input: Set feasible set $\Theta$, initial control parameters $\Theta_{\text{init}}$, weight vector $\mathrm{w}$, DT sampling length $\tilde{n}$, DT activation parameters $\eta_{1}$ and $\eta_{2}$, DT re-initialization parameter $\tilde{\delta}$, maximum experiments $n_{\text{max}}$, DT stopping threshold and consecutive iteration parameters $\eta_{3}$ and $\tilde{n}_{\text{EI}}$
@@
3:while $n_{\text{max}}$ experiments on $\mathrm{G}$ is not reached do
@@
7:  if DT activation criteria in (12) is met then
@@
[…]
@@
16:   Set $\theta^{+}=\theta_{m}\in\mathcal{D}\text{ s.t. }\forall\theta\in\mathcal{D}:J(\theta_{m})<J(\theta)$
[…]
@@
23:  if DT re-initialization criterion in (13) is met then
@@
An appropriate stopping criterion for the optimization process is necessary for the practical implementation of the algorithm. We terminate the BO after a fixed number of iterations on the real system denoted by $n_{\max}$. When using DT, we need a stopping condition that ensures prompt termination of the optimization. We use the stopping criterion that depends on the maximal expected improvement over previous iterations and the current expected improvement proposed in [35]. Once we have activated the DT plant to predict the closed-loop response, we stop the BO iterations on DT when the maximum number of iterations on the DT is reached or depending on the ratio between the last expected improvement and the maximal expected improvement realized so far. This way, we avoid unnecessary iterations using the DT plant where we anticipate no further improvement. Notably, we stop BO on DT when for $\tilde{n}_{{\text{EI}}}$ consecutive iterations on DT we meet one of the conditions
$a_{\text{EI},i}(\theta)\ \displaystyle\leq$ $\ \eta_{3}\>\underset{i-\tilde{n}_{EI}\leq j\leq i-1}{\mathrm{max}}\>a_{\text{{\text{EI}}},\text{j}}(\theta),$ (14a)
$i\$$>$ $\ n_{\text{max}},$ (14b)
where $i$ is the number of BO iterations on DT after last time DT is activated, $a_{\text{EI},i}(\theta)=a_{\text{EI}}(\theta^{+})$ and $\eta_{3}>0$ is a predefined threshold to stop BO iteration on DT as in [50]. If the expected improvement does not significantly improve compared with the improvements in the last $\tilde{n}_{EI}$ iterations, then (14a) halts the optimization with DT.
[…]
The training data set size $n$ changes when we remove the DT data from $\mathcal{D}$ (see line 17 of Algorithm 1). Furthermore, at each iteration $m$ on the real plant (where $0\leq m\leq n_{\text{max}}$), the DT can be activated and $i$ number of BO iterations can be performed on the DT where $i\leq n_{\text{max}}$. So the computational complexity of the guided BO algorithm is bounded and dominated by a polynomial that depends on $i$, $m$, and $n$.
@@ IV-A System Structure and Identification
We measure the plant response to a given sinusoidal signal with various frequencies, a fixed amplitude, and a single DC value of $\bar{\Omega}=5522.40\ \text{rad/s}$. Considering several excitation frequencies, the excitation signal has a fixed amplitude of 613.60rad/s613.60\ $\text${rad/s}. We measure the system output with a high sampling rate of $0.5$ ms and draw the bode diagram of the system response, which allows us to fit a second-order LTI transfer function visually. We repeat each measurement two times to alleviate the measurement noise. For A/D conversion, we sample 128 points evenly distributed in every period, introducing a constant phase shift in our bode diagram. So, because we have $12$ different frequencies for the sinusoidal excitation signal, the sampling rate varies (see Table I). Consequently, an input time delay of $2$ ms abstracts the time shift introduced by A/D conversion. We address this delay by shifting the phase in the bode diagram to a fixed offset corresponding to the delay. Finally, $\mathrm{G}$ is written as $\mathrm{G}(s):=\frac{L}{s^{2}+\tau_{1}s+\tau_{2}}\,\mathrm{e}^{-sl}$, where Table I summarizes its identified parameters.